OpenStudy (anonymous):

Find the distance between (0, 0) and (5, 8). Round the answer to the nearest hundredth.

OpenStudy (zzr0ck3r):

The distance between \((a,b)\) and \((c,d)\) is given by \[d=\sqrt{(a-c)^2+(b-d)^2}\]

OpenStudy (anonymous):

yeah i got that but my answer isnt one of the options

OpenStudy (anonymous):

a. 9.68 units c. 9.43 units b. 10.38 units d. 8.75 units

OpenStudy (anonymous):

8.54400374531753

OpenStudy (zzr0ck3r):

\(\sqrt{(5-0)^2+(8-0)^2}=\sqrt{5^2+8^2}=\sqrt{64+25}=\sqrt{89}\approx 9.43398\)